Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Wavefront sensing)
chiummo, rocchi, menzione - 23:32 Wednesday 22 March 2023 (59385) Print this report
HWS-INJ SLED beam alignment

Goal of the shift was to re-align the SLED beam on the West ITM.

Antonino went to the EIB, blocked the bench, and Nicola locked the West arm on the green.

Antonino re-aligned the SLED beam on the green beam, acting on POL1 and EIB_Har_M0. We then noticed that the beam was mis-aligned on the sensor, so he moved EIB_Har_M5 to re-center it, with the ITF now in single bounce. As a further check of the good alignment of the SLED beam, we asked Nicola to re-align the WE to see if there the effect of the reflection from the ETM on the HWS image: indeed we could see a significant increase of the CCD counts, as expected localized in a small portion of the HWS image.

To check the status of the alignment, we decided to start a HWS-INJ acquisition to see the OPL induced by the CH.

At 18:28 UTC, HWS-INJ acquisition started (/data/tcs/buffer/HWS-RC/INJ/20230322T1928)

At 18:39 UTC, WI CH flipper closed, to see the cooldown curvature.

At 18:55 UTC, since nothing was appearing on the HWS maps, I decided to open the CH flipper mirror and increse the power by 300 mW

19:15 UTC, my ThinLinc session crashed, and HWS acquisition stopped.

19:18 UTC, at this point I decided to start over, so I closed the WI CH flipper and wait for the cool down.

20:28 UTC, HWS-INJ acquisition started (/data/tcs/buffer/HWS-RC/INJ/20230322T2128)

20:33 UTC, WI CH flipper opened (pickoff from 270 mW to 2370 mW)

21:00 UTC, comparing the HWS maps before (fig. 1) and after (fig. 2) switching on the CH, the beam seems to be on the lower left corner of the image.

21:02 UTC, HWS-INJ acquisition stopped.

21:07 UTC, standard CH power restored (65 mW).

The CCD illumination pattern has to be re-checked.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×